OpenStudy (anonymous):

int ( 2x / ( 1 - cos 2x) )dx ; anybody ?

OpenStudy (anonymous):

help needed guys . \[\int\limits (2x / ( 1+ \cos 2x)) dx\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

lol let lalaly finish

OpenStudy (anonymous):

@nickhouraney ; okay and @lalaly : no pressure take your time .

OpenStudy (lalaly):

\[\int\limits{\frac{2x}{1+\cos(2x)}dx}\]sincee\[\frac{1}{2}(1+\cos(2x))=\cos^2x\]so integration becomes\[\frac{1}{2}\times 2 \int\limits{\frac{x}{\frac{1}{2}(1+\cos(2x))}}dx\] \[=\int\limits{\frac{x}{\cos^2x}}dx\]\[=\int\limits{xsec^2x}dx\] now solve by parts u=x dv=sec^2xdx du=dx v=tanx so integral becomes\[xtan(x)-\int\limits{\tan(x)dx}\] do u know how to integrate tan(x)

OpenStudy (anonymous):

Forgot

OpenStudy (anonymous):

Sorry

OpenStudy (lalaly):

ok then\[\tan(x)=\frac{\sin(x)}{\cos(x)}\] so\[\int\limits{\tan(x)}dx=\int\limits{\frac{\sin(x)}{\cos(x)}}dx\]now let u=cosx du=-sin(x)dx substitute in the integral\[\int\limits{\frac{-du}{u}}\]\[=-\ln(u)\]substitute u=cos(x)\[\int\limits{\tan(x)dx}=-\ln(\cos(x))+C\] so\[\int\limits{\frac{2x}{(1+\cos(2x)}}dx=xtan(x)-\ln(\cos(x))+c\]

OpenStudy (anonymous):

Thank you very much ; miss high school .

OpenStudy (lalaly):

ur welcome